OpenStudy (watermelon14):

The table below compares the radioactive decay rates of two materials. Material Original mass of material (in grams) Mass of material after 60 hours (in grams) 1 208 13 2 200 50 Based on the table, which of these conclusions is most likely correct? The half-life of Material 1 and Material 2 are equal. The half-life of Material 2 is double the half-life of Material 1. The half-life of Material 2 is 30 hours more than the half-life of Material 1. The half-life of Material 1 is 30 hours more than the half-life of Material 2.

OpenStudy (festinger):

In 60 hours, sample 1 went through 4 half lives (13/208 = 1/16 = 1/2 * 1/2 * 1/2* 1/2) In 60 horus, sample 2 went through 2 half lives (50/200=1/4 = 1/2 * 1/2) Thus, sample 1 has a half life of 60/4 = 15h and sample 2 has a half life of 60/2 = 30h so answer is: The half-life of Material 2 is double the half-life of Material 1.

OpenStudy (watermelon14):

Thank you so very much! :P